An audacious physics student weighing 130 pounds a
Answer #1

The free body diagram is:

Fr Girl

And the sum of forces when the string is streched is:

\sum F = F_{R} - W = 0

F_{R} = W

-kx = mg

x = \frac{mg}{k}

x = \frac{(58.97kg)(9.8m/s^{2})}{540N/m}

x = 1.07m = 32.61ft

This is the length that the string streched. Now, we sum this with the length of the string unstreched to see if it passes the value of the fall:

L = 110ft + 32.61ft

L = 142.61ft

ANSWER: she makes it by a little. The important thing to see here is that the string is almost the length of the fall, and to do this it is important to know how is the maximum streching to avoid accidents.

I hope it helps!!